# Rethinking Pan-sharpening: Principled Design, Unified Training, and a Universal Loss Surpass Brute-Force Scaling

### v2

@@ 3.3 Loss Function
4 Experiments
4.1 Setup
Datasets. We conduct experiments on three public datasets: WorldView-2 (WV2), WorldView-3 (WV3), and GaoFen-2 (GF2). For our primary “all-in-one” experiments, we combine the training sets of all three. We follow standard protocols for evaluation, using both reduced-resolution and full-resolution test sets.
Evaluation Metrics. We provide a comprehensive evaluation using both reference and no-reference metrics. For reduced-resolution evaluation, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM) (Wang et al. 2004), Spectral Angle Mapper (SAM) (Yuhas, Goetz, and Boardman 1992), and ERGAS (Wald 2002). For full-resolution evaluation, we use the no-reference metrics Dλ, Ds, and QNR (Alparone et al. 2008).
Implementation Details. Our framework is implemented in PyTorch. All models are trained on a single NVIDIA RTX 4090 GPU. We use the ADAM optimizer with a learning rate of $5\times 10^{-4}$ and betas of (0.9,0.999)(0.9,0.999)( 0.9 , 0.999 ). A cosine annealing scheduler adjusts the learning rate over 500 epochs with a batch size of 16.
4.2 Quantitative Comparison and Ablations
Our experimental evaluation is designed to validate two core theses: 1) our “all-in-one” training paradigm is a more robust and effective method for developing generalizable pan-sharpening models, and 2) our “PanTiny” architecture achieves a superior balance of performance and efficiency compared to existing methods.
Model Training WV2 WV3 GF2
[…]
Pan-Mamba (He et al. 2025) All-in-One 41.39 0.9663 0.0236 30.17 0.9174 0.0779 43.98 0.9725 0.0164
Separate 42.24 0.9729 0.0212 31.16 0.9299 0.0702 47.65 0.9894 0.0103
PNN (Masi et al. 2016) All-in-One 39.82 0.9540 0.0282 29.49 0.9005 0.0861 43.14 0.9667 0.0178
Separate 40.76 0.9624 0.0259 29.94 0.9121 0.0824 43.12 0.9704 0.0172
PanFlow (Yang et al. 2023) All-in-One 41.11 0.9645 0.0243 30.04 0.9106 0.0799 46.36 0.9825 0.0125
Separate 41.86 0.9712 0.0224 30.49 0.9221 0.0751 47.25 0.9884 0.0103
PSCINN (Wang et al. 2024) All-in-One 35.60 0.8967 0.0336 22.61 0.5538 0.1115 42.69 0.9616 0.0181
Separate 41.85 0.9703 0.0223 30.56 0.9230 0.0748 47.11 0.9878 0.0107
CFDCNet (Li et al. 2025) All-in-One 41.54 0.9667 0.0233 30.42 0.9155 0.0775 47.76 0.9866 0.0107
Separate 42.24 0.9733 0.0209 31.24 0.9327 0.0694 47.84 0.9902 0.0097
Pan-LUT (Cai et al. 2025) All-in-One - - - - - - - - -
Separate 39.84 0.9555 0.0286 28.82 0.8936 0.0935 42.66 0.9642 0.0189
Separate 42.16 0.9711 0.0217 30.61 0.9245 0.0747 48.93 0.9900 0.0092
Main Results on Multi-Dataset Training
Table 1 presents the main results of our study. All listed methods were trained under our unified “all-in-one” paradigm on the combined WV2, WV3, and GF2 datasets, and evaluated on each one’s test set using a single set of model weights. We present two versions of our model: PanTiny (Small), our ultra-lightweight variant, and PanTiny (Big), our primary model that achieves the best performance. We include both to highlight the excellent efficiency of our base architecture and the SOTA performance achieved with a modest increase in size.
Our proposed PanTiny (Big) achieves SOTA performance across all three datasets, outperforming both classic and recent methods. Notably, it surpasses CFDCNet (Li et al. 2025), a much larger model, on all metrics. It also significantly outperforms other lightweight methods like PanFlow (Yang et al. 2023). The results for PSCINN (Wang et al. 2024) highlight their instability in a multi-dataset setting, as it failed to complete training, further validating our design choices for robustness.
Impact of the All-in-One Training Paradigm
A core contribution of our work is the “all-in-one” training paradigm. Prior work on generalization often involves training on a single dataset and testing on others. However, due to the significant domain gap between satellite datasets, this approach struggles to produce a truly universal model. Our preliminary tests (detailed in the appendix) show that a model trained for just one epoch on a source dataset can sometimes match the cross-dataset performance of a fully-trained so-called “general” model, suggesting the latter may be overfitting.
Our “all-in-one” approach directly addresses this by exposing the model to multiple domains during training. As shown in Table 3, this has a profound effect. Complex models like Pan-Mamba and PSCINN see a significant performance drop compared to their specialized, separately trained counterparts. In contrast, our “PanTiny (Big)” model is remarkably robust, with only a minor drop of  0.3 PSNR on WV2/GF2 and almost no change on WV3.
Furthermore, Table 4 demonstrates that this paradigm enhances generalization on real-world, full-resolution data. For all tested models, switching from separate to all-in-one training results in a substantial improvement in the no-reference QNR metric on the WV2 full-resolution dataset. While PanFlow achieves the highest QNR, the universal improvement underscores the power of our proposed training method as a general tool for improving pan-sharpening model generalization. Note that some models like Pan-Mamba and CFDCNet are excluded as their codebase does not support variable inference resolutions without modification; details are in the supplement.
Model QNR $\uparrow$
Separate All-in-One
MSDCNN (Yuan et al. 2018) 0.7683 0.8898
PanNet (Yang et al. 2017) 0.7684 0.8726
PNN (Masi et al. 2016) 0.7527 0.8844
PSCINN (Wang et al. 2024) 0.7904 0.8849
PanTiny (Small) 0.7827 0.8751
PanTiny (Big) 0.7985 0.8793
PanFlow (Yang et al. 2023) 0.7910 0.8900
Model Params(K) FLOPs(G) WV2 WV3 GF2
[…]
CFDCNet (Li et al. 2025) 1700.8 55.73 42.50 0.9729 0.0205 31.11 0.9298 0.0707 49.07 0.9903 0.0091
Pan-Mamba (He et al. 2025) 488.8 16.02 41.77 0.9691 0.0226 30.36 0.9215 0.0769 45.84 0.9811 0.0134
PanFlow (Yang et al. 2023) 87.3 2.86 41.68 0.9688 0.0229 30.24 0.9197 0.0785 47.49 0.9865 0.0109
MSDCNN (Yuan et al. 2018) 239.0 7.83 41.46 0.9669 0.0236 30.18 0.9189 0.0786 44.09 0.9730 0.0161
PNN (Masi et al. 2016) 68.9 2.26 40.84 0.9635 0.0256 29.82 0.9128 0.0834 43.40 0.9688 0.0173
PanNet (Yang et al. 2017) 80.3 2.63 40.79 0.9620 0.0256 29.82 0.9106 0.0841 43.76 0.9705 0.0167
Model Params WV2 WV3 GF2
(K) PSNR SSIM PSNR SSIM PSNR SSIM
PanTiny (Small) 48.3 41.62 0.9685 30.38 0.9216 48.16 0.9884
PanTiny (Big) 81.7 41.85 0.9696 30.59 0.9238 48.61 0.9894
PanTiny (Large Body) 172.4 42.06 0.9708 30.67 0.9248 48.75 0.9896
PanTiny (Huge Body) 195.9 42.12 0.9711 30.74 0.9258 48.85 0.9898
Architectural Design and Efficiency
To demonstrate the superiority of our network architecture, we conducted a benchmark where all competing methods were trained with our proposed composite loss. As shown in Table 5, even when other methods benefit from our improved training process, our “PanTiny (Big)” model still secures a top-tier position, surpassed only by the much larger CFDCNet (+1.6M params, +53G FLOPs). Our exploration process, also detailed in the table, further justifies our choices:
  • Limitations of Naive Scaling: We first explored two intuitive designs: “DeepPNN”, a deeper and wider version of PNN, and “ResAtten”, which combines a standard ResNet backbone with attention. While both achieve competitive performance (with ResAtten reaching the highest PSNR on WV2), they require a significantly larger number of parameters (over 260K). This demonstrates that simply scaling up or using generic vision backbones is not the most efficient path to SOTA performance.
  • Efficiency of PanTiny (Small): Our purpose-built single-encoder model achieves strong results across the board while being one of the smallest models we are aware of in the literature, with only 48.3K parameters—even smaller than other lightweight methods like SFDI.
The effect of scaling is further explored in Table 6. By expanding our model to a ‘Huge’ version (196K params), we can nearly match the performance of CFDCNet on GF2 (48.85 vs 49.07 PSNR), but at less than 12% of its parameter count. This reinforces our core argument against inefficient scaling and validates our choice of “PanTiny (Big)” as the optimal model.
Fusion and Refinement Modules
We ablate the fusion and refinement blocks in Table 7 and Table 8. For fusion, our simple “Enhanced Conv” (2-layer Conv) outperforms both a simpler “$\times 1$ 1 Conv” and more complex attention-based mechanisms. Notably, “DeepFusion” from Pan-Mamba, which is highly effective in a single-dataset setting, performs poorly here, again suggesting it overfits. For refinement, a simple “Conv” layer is optimal. Adding complexity via a larger convolution or channel attention increases parameters without a consistent performance benefit, and in some cases, hurts the results.
@@ 4.2 Quantitative Comparison and Ablations
Loss Function
The significant performance boost of our method stems not only from its architecture but also from our carefully engineered loss function—a key innovation in its own right. As shown in Table 5, our composite loss provides a universal performance uplift to all tested models, demonstrating its power as a general tool for the community. Our extensive search for the optimal configuration, detailed in the appendixs, began with evaluating individual components. This revealed that SSIM loss is particularly effective, especially for the GF2 dataset. Building on this insight, we found that assigning a high weight to the SSIM component consistently yielded improvements. Our final weights ($\lambda_{1}=1.5,\lambda_{2}=4.0,\lambda_{3}=1.5$) represent the best-performing combination from this exhaustive search. This well-tuned, composite loss has proven to be a cornerstone of our work, enabling a significant leap in performance and pushing the pan-sharpening field into a new era of 48-49 PSNR on the GF2 dataset.
@@ 5 Conclusion
In this paper, we challenged the prevailing “bigger is better” paradigm in pan-sharpening by focusing on efficiency, generalization, and principled engineering. We introduced PanTiny, a lightweight and highly efficient model, and demonstrated that its carefully considered architecture achieves a superior balance of performance and computational cost compared to larger, more complex SOTA models. Our most significant contribution is the novel “all-in-one” training paradigm. By training a single model jointly on multiple diverse datasets, we not only simplified the deployment pipeline but also demonstrably improved the generalization capabilities of various models on real-world, full-resolution data. This stands in contrast to previous generalization efforts, which often struggle with domain gaps.Finally, we presented a universally effective composite loss function that provides a significant performance uplift across all tested architectures, pushing the benchmarks for the field into a new era. We believe that our combined contributions—the PanTiny model, the all-in-one paradigm, and the powerful loss function—offer a more sustainable and practical path forward for future pan-sharpening research. Our code and pre-trained models will be made publicly available in the appendixs.
Appendix A On the Limitations of Generalization and the Necessity of All-in-One Training
@@
A.1 The Challenge of Domain Gaps in Pan-sharpening Datasets
@@
A.2 The “One-Epoch Generalization” Illusion
To test the hypothesis of overfitting in conventional generalization studies, we conducted a surprising experiment: we trained several of our intermediate models for only one epoch on the WV2 dataset and then evaluated their performance on the unseen WV3 and GF2 datasets. The results, shown in Table 10, are striking. When tested on GF2, our one-epoch trained “M4 (Channel Attn)” model achieves a PSNR of 39.13. This result is comparable to or even surpasses the performance of fully-trained models from dedicated generalization papers, such as DDIF (Chen et al. 2022), which reports a PSNR of 37.77 on GF2 after being fully trained on WV2 (see Table 9). This suggests that the hundreds of additional training epochs in those works contribute little to true generalization, instead primarily reinforcing the model’s bias towards the source dataset. This finding strongly motivates a shift away from the “train-on-one, test-on-many” methodology. Any reviewer can easily verify this conclusion with a personal computer in under 5 minutes using our provided codebase, if they already have the datasets.
@@ A.2 The “One-Epoch Generalization” Illusion
Table 9: Quantitative comparison from a prior generalization work (Chen et al. 2022), with the model trained on the Worldview-II dataset and tested on other datasets. The best results are marked in bold and the second results are marked with underline. $\uparrow$ indicates that the larger the value, the better the performance, and $\downarrow$ indicates that the smaller the value, the better the performance.
@@
A.3 The Overfitting Trap of Separate Training
@@
A.4 Failure Case: Generalization to Jilin-1 Dataset
@@
Appendix B Detailed Ablation on the Composite Loss
B.1 The Overlooked Potential of Loss Functions
Historically, the pan-sharpening community has predominantly focused on advancing model architectures to achieve performance gains. The L1 loss has long been the de-facto standard, with the majority of research efforts dedicated to designing more sophisticated networks. However, this model-centric approach appears to be reaching a point of diminishing returns. As evidenced by recent SOTA models like CFDCNet (Li et al. 2025), achieving marginal performance improvements now requires an enormous increase in computational cost (over 55G FLOPs), suggesting an architectural bottleneck.
We posit that the loss function, a relatively underexplored area, holds the key to unlocking the next level of performance. While perceptual losses like SSIM (Wang et al. 2004) have been considered, they were often dismissed after preliminary tests showed that using them in isolation or with balanced weights did not yield superior results and could sometimes introduce color artifacts. This led to a widespread underestimation of their potential. We believe that a systematic, large-scale exploration of loss combinations has been a missing piece in the field.
B.2 Our Systematic Two-Stage Search for the Optimal Loss
@@
Appendix C Detailed Ablation on Model Architecture
@@ C.1 Downsampling Strategy
C.2 Investigating the “DeepFusion” Module
@@
C.3 Single-Encoder vs. Dual-Encoder Design
@@
C.4 Full Ablation Results for Final Model Components
@@
Appendix D Additional Visual Results

### v3

@@ 4.3.2 Architectural Design and Efficiency
4.4 Ablation Studies and Further Analysis
We validate the specific design choices for our loss function and the PanTiny architecture.
4.4.1 PanTiny’s Architectural Ablations
We now validate the internal design choices of PanTiny to demonstrate that its simplicity is a principled and effective strategy. We conducted detailed ablations on its core fusion and refinement components.
Fusion Module: The fusion mechanism is critical for integrating PAN and MS information. As shown in Table 7, we tested several alternatives. A simple “1x1 Conv” provides a strong baseline, but our proposed “Enhanced Conv” (a 2-layer Conv) achieves the best overall performance. Notably, we also tested the complex “DeepFusion” module from Pan-Mamba [8]. This module, while effective in its original single-dataset setting, performs poorly here. This result strongly suggests that complex, specialized modules tend to overfit to dataset-specific artifacts, making them less suitable for the robust, multi-dataset ”all-in-one” paradigm. This validates our choice of a simpler, more robust block.
Refinement Module: We applied the same principle to the refinement block, as shown in Table 8. Our final model uses a simple “Conv” layer. We found that adding complexity, such as a “Large Conv” or a “Channel Attn.” block, unnecessarily increases the parameter count without yielding a consistent or significant performance benefit. These ablations on both the fusion and refinement modules confirm that PanTiny’s simplicity is not naive, but rather a deliberate and empirically-validated design choice that prioritizes robustness and parameter efficiency in our unified training framework.
@@ 4.4.1 PanTiny’s Architectural Ablations
4.4.2 Universality of the Composite Loss
To demonstrate the superiority of our loss function as a general tool, we conducted a benchmark where all competing methods were trained with it (Table 4). The pan-sharpening community has historically relied heavily on the standard L1 loss, often used in isolation, as it is simple and stable. However, our initial ablations in Table 6 show that using a single loss component, including L1, yields suboptimal performance.
This motivated an extensive, systematic loss search to find an optimal composite weighting. Our final weights ($\lambda_{1}=1.5,\lambda_{2}=4.0,\lambda_{3}=1.5$) are the result of this search, which is detailed further in our supplementary material. A key insight from this process was that while L1 provides a stable regression base, assigning a particularly high weight to the SSIM component ($\lambda_{2}=4.0$) consistently delivered the best performance across datasets, especially for GF2. As shown in Table 4, when competing models (often trained with a simple L1 loss in a ‘Separate’ paradigm) are retrained with our unified strategy and this optimized composite loss, they see a significant performance uplift. This demonstrates our composite loss’s power as a general tool for the community.
@@ 5 Conclusion
In this paper, we challenged the pan-sharpening community’s prevailing focus on model-centric innovation and complex cross-dataset generalization. We shifted the focus to two more practical, yet overlooked, challenges: full-resolution (FR) generalization and deployment usability. Our primary contribution is the novel “all-in-one” training paradigm, which we proved is a universally effective strategy for dramatically boosting the FR generalization (e.g., QNR) of all tested models. This paradigm also inherently solves the “one-model-per-dataset” problem, and we support this with a highly reproducible, dependency-free codebase for true usability. Finally, we proposed PanTiny, a lightweight model designed for this new paradigm. We demonstrated that this simple, efficient architecture is the SOTA performer in this practical setting, proving that principled, robust design is more effective than the complex architectural scaling seen in other works. We believe our combined contributions—the paradigm for FR generalization, the focus on usability, and the efficient PanTiny model and a universal composite loss—offer a more sustainable and practical path forward for pan-sharpening research.
Supplementary Material
6 On the Limitations of Generalization and the Necessity of All-in-One Training
@@
6.1 The Challenge of Domain Gaps in Pan-sharpening Datasets
@@
6.2 The “One-Epoch Generalization” Illusion
To test the hypothesis of overfitting in conventional generalization studies, we conducted a surprising experiment: we trained several of our intermediate models for only one epoch on the WV2 dataset and then evaluated their performance on the unseen WV3 and GF2 datasets. The results, shown in Table 10, are striking. When tested on GF2, our one-epoch trained “M4 (Channel Attn)” model achieves a PSNR of 39.13. This result is comparable to or even surpasses the performance of fully-trained models from dedicated generalization papers, such as DDIF [6], which reports a PSNR of 37.77 on GF2 after being fully trained on WV2 (see Table 9). This suggests that the hundreds of additional training epochs in those works contribute little to true generalization, instead primarily reinforcing the model’s bias towards the source dataset. This finding strongly motivates a shift away from the “train-on-one, test-on-many” methodology. Any reviewer can easily verify this conclusion with a personal computer in under 5 minutes using our provided codebase, if they already have the datasets.
@@ 6.2 The “One-Epoch Generalization” Illusion
Table 9: Quantitative comparison from a prior generalization work [6], with the model trained on the Worldview-II dataset and tested on other datasets. The best results are marked in bold and the second results are marked with underline. $\uparrow$ indicates that the larger the value, the better the performance, and $\downarrow$ indicates that the smaller the value, the better the performance.
@@
6.3 The Overfitting Trap of Separate Training
@@
6.4 Failure Case: Generalization to Jilin-1 Dataset
@@
7 Detailed Ablation on the Composite Loss
7.1 The Overlooked Potential of Loss Functions
Historically, the pan-sharpening community has predominantly focused on advancing model architectures to achieve performance gains. The L1 loss has long been the de-facto standard, with the majority of research efforts dedicated to designing more sophisticated networks. However, this model-centric approach appears to be reaching a point of diminishing returns. As evidenced by recent SOTA models like CFDCNet [11], achieving marginal performance improvements now requires an enormous increase in computational cost (over 55G FLOPs), suggesting an architectural bottleneck.
We posit that the loss function, a relatively underexplored area, holds the key to unlocking the next level of performance. While perceptual losses like SSIM [16] have been considered, they were often dismissed after preliminary tests showed that using them in isolation or with balanced weights did not yield superior results and could sometimes introduce color artifacts. This led to a widespread underestimation of their potential. We believe that a systematic, large-scale exploration of loss combinations has been a missing piece in the field.
7.2 Our Systematic Two-Stage Search for the Optimal Loss
@@
8 Detailed Ablation on Model Architecture
@@ 8.1 Downsampling Strategy
8.2 Investigating the “DeepFusion” Module
@@
8.3 Single-Encoder vs. Dual-Encoder Design
@@
8.4 Full Ablation Results for Final Model Components
@@
9 Additional Visual Results